OpenStudy (anonymous):

at the given point fine the slope of the curve the line that is tangent to the curve or the line that is normal to the curve. x^3y^3=8 tangent at (2,1)

OpenStudy (rogue):

x^3 y^3 = 8 xy = 2 y = 2/x y' = -2/x^2 y' (2) = -2 / 4 = -1/2 Ytangent = (-1/2)(x - 2) + 1